what is the average speed of a bus that increases uniformly from 15m/s to 60m/s in 20 seconds


1
Expert's answer
2021-04-26T17:09:08-0400
"v_{av}=0.5(v_i+v_o)\\v_{av}=0.5(15+60)=37.5\frac{m}{s}"